# EIDSeg: A Pixel-Level Semantic Segmentation Dataset for Post-Earthquake Damage Assessment from Social Media Images

### v1

@@ Conclusion
We introduce EIDSeg, the largest pixel-level semantic segmentation dataset for post-earthquake damage analysis based on social media images. The dataset comprises 3,266 images collected from nine major earthquakes that occurred between 2008 and 2023, covering a diverse range of real-world disaster scenarios. Annotations span five semantic classes: Undamaged Building, Damaged Building, Destroyed Building, Undamaged Road, and Damaged Road, allowing a fine-grained evaluation of damage to built infrastructure. Through a carefully designed annotation guideline and expert-led quality control, EIDSeg provides a reliable foundation for training vision models capable of automating post-earthquake damage assessment using ground-view imagery.
@@ Acknowledgment
The research described herein was supported in part by the Elizabeth and Bill Higginbotham Professorship at Georgia Tech and the School of Computing Instruction at Georgia Tech. Special thanks to Mohammad Taher for helping with the annotation. The supports are gratefully acknowledged.
Appendix
A. Annotators Misalignment Examples
@@
  • Pancake damage infrastructure (R1). The image shows a partially collapsed building where upper floors have failed while lower floors remain standing. Annotation A labels the entire structure as a damaged building, while Annotation B assigns the collapsed portion to Debris. Following DSS annotation practice (?), we segment pancake collapsed components as Debris rather than Damaged Building.
@@ C. Histogram of the Images before Data Cleaning
Figure 7: Distribution of image widths and heights in the EID dataset (?) The red dashed lines indicate the mean values, and the orange dashed lines indicate the medians. Skewness and kurtosis are reported above each histogram, highlighting the strong positive skew and heavy-tailed nature of the resolution distribution. The $x$-axis is truncated at 2000 pixels for clarity, since images with larger dimensions are rare.
@@ D. Leave-One-Event-Out Results
The leave-one-disaster-out evaluation with the EoMT model requires identifying the source disaster for every image. Because EID (?) and DSS (?) aggregate material from multiple online sources without full provenance, the analysis includes only images with verifiable origins. This restriction ensures that every split corresponds uniquely to one event.

### v2

@@ Conclusion
In this work, we introduce EIDSeg, the largest pixel-level semantic segmentation dataset for post-earthquake damage analysis based on social media images. The dataset comprises 3,266 images collected from nine major earthquakes that occurred between 2008 and 2023, covering a diverse range of real-world disaster scenarios. Annotations span five semantic classes: Undamaged Building, Damaged Building, Destroyed Building, Undamaged Road, and Damaged Road, allowing a fine-grained evaluation of damage to built infrastructure. Through a carefully designed annotation guideline and expert-led quality control, EIDSeg provides a reliable foundation for training vision models capable of automating post-earthquake damage assessment using ground-view imagery.
@@ Acknowledgment
The research described herein was supported in part by the School of Computing Instruction and the Elizabeth and Bill Higginbotham Professorship at Georgia Tech. This support is gratefully acknowledged. Special thanks are due to Mohammad Taher for helping with the annotation. Support for the undergraduate students participating in the project was provided by the US National Science Foundation through the Geotechnical Engineering Program under Grant No. CMMI-1826118. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of NSF.
Appendix A Appendix
@@ A. Annotators Misalignment Examples
  • Pancake damage infrastructure (R1). The image shows a partially collapsed building where upper floors have failed while lower floors remain standing. Annotation A labels the entire structure as a damaged building, while Annotation B assigns the collapsed portion to Debris. Following DSS annotation practice (Zhang2025PixelsDamageSeverity), we segment pancake collapsed components as Debris rather than Damaged Building.
@@ C. Histogram of the Images before Data Cleaning
Figure 7: Distribution of image widths and heights in the EID dataset (Huang2025Fidelity) The red dashed lines indicate the mean values, and the orange dashed lines indicate the medians. Skewness and kurtosis are reported above each histogram, highlighting the strong positive skew and heavy-tailed nature of the resolution distribution. The $x$-axis is truncated at 2000 pixels for clarity, since images with larger dimensions are rare.
@@ D. Leave-One-Event-Out Results
The leave-one-disaster-out evaluation with the EoMT model requires identifying the source disaster for every image. Because EID (Huang2025Fidelity) and DSS (Zhang2025PixelsDamageSeverity) aggregate material from multiple online sources without full provenance, the analysis includes only images with verifiable origins. This restriction ensures that every split corresponds uniquely to one event.